OpenStudy (anonymous):

Find the replacement 15 - x = 9 A) 5 B) 6 C) 7 D) 8

OpenStudy (triciaal):

what do you need to do to isolate x to one side of the equation?

OpenStudy (anonymous):

Answer 6 @Fishy1237

OpenStudy (midhun.madhu1987):



OpenStudy (anonymous):

Why don't you simply try each number as a replacement for x? Find the replacement 15 - x = 9 A) 5 B) 6 C) 7 D) 8 15 - 5 = 9 ? Wrong, 15-5 = 10 15 - 8 = 9 ? Wrong, 15-8 = 7 Your fastest way of answering the question is to pick the right answer, the one that makes 15 - ___ = 9 What did you get?